4x^3-54x^2+180x-125=0
: 4x^3-54x^2+180x-125=0

Answer by mangopeeler07(445)

4x^3-54x^2+180x-125=0

Group terms
(4x^3-54x^2)+(180x-125)=0

Factor
2x^2(2x-27)+5(36x-25)=0

Set each expression equal to 0
2x^2=0
x^2=0
x=0

2x-27=0
2x=27
x=27/2

36x-25=0
36x=25
x=25/36

Answer:
x=0 or 27/2 and 25/36